OpenStudy (anonymous):

perform the indicated operation and simplify if possible - 24/55 / 39/40

sam (.sam.):

\[\Huge \frac{(\frac{-24}{55})}{(\frac{39}{40})}\]?

OpenStudy (anonymous):

its the negative sign and then the 24/55 so im guessing that both numbers must be negative or something

sam (.sam.):

\[\Huge -\frac{(\frac{24}{55})}{(\frac{39}{40})}?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

the negative sign is beside the line in between the 24 and 55 its not beside the 24

sam (.sam.):

its same

OpenStudy (anonymous):

ok so I never knew that sorry

sam (.sam.):

\[\Huge -\frac{(\frac{24}{55})}{(\frac{39}{40})}=-\frac{24}{55} \div \frac{39}{40}\] \[\large -\frac{24}{55} \div \frac{39}{40}\] Switch the divide to multiply, which makes one of the fraction switch \[\large -\frac{24}{55} \times \frac{40}{39}\] Then multiply them together, numerator with numerator, denominator with denominator